# Diffusion Model as a Noise-Aware Latent Reward Model for Step-Level Preference Optimization

### v2

@@ 4.1 Latent Reward Model
As described in Sec. 1, the pre-trained diffusion model is remarkably suitable for step-level reward modeling due to its inherent ability to align texts and images, extract features directly from $x_{t}$ at any timestep, and naturally accept timestep inputs. Therefore, we introduce the Latent Reward Model (LRM) to employ diffusion models for reward modeling.
Architecture of LRM. LRM leverages features of the UNet and text encoder in the diffusion model for preference prediction, as depicted in Fig. 4. Specifically, the textual features $f_{p}\in\mathbb{R}^{n_{l}\times n_{p}}$ are extracted from the prompt $p$ by the text encoder, where $n_{l}$ and $n_{p}$ denotes the number of tokens and the dimension of textual features, respectively. Following CLIP (Radford et al., 2021), we use the last token $f_{eos}\in\mathbb{R}^{1\times n_{p}}$ to represent the entire prompt and incorporate a text projection layer to obtain the final textual features $T\in\mathbb{R}^{1\times n_{d}}$, where $n_{d}$ represents the final dimension of textual and visual features. Each noisy latent image $x_{t}$ is passed through the UNet to interact with textual features $f_{p}$. The visual features of UNet are averagely pooled along the spatial dimension, resulting in the multiscale down-block features $V_{down}$ and middle-block features $V_{mid}$ as follows
$V_{down},V_{mid}=\text{AvgPool}(\text{UNet}(x_{t},f_{p})),$ (8)
where $V_{down}=\{V_{d_{i}},i=1,\ldots,L\}$. $L$ is the number of down blocks and $V_{d_{i}}$ represents features of the $i$-th down blocks. However, it is observed that these features lack sufficient correlations with the textual features. Inspired by the Classifier-Free Guidance (Ho & Salimans, 2022), we propose the Visual Feature Enhancement (VFE) module to enhance correlations between visual and textual features. It first extracts middle-block features $V_{mid\_ucond}$ that are independent of textual information following a process similar to extracting $V_{mid}$, but with a null prompt. Then the enhanced visual features $V_{enh}\in\mathbb{R}^{1\times n_{m}}$ can be calculated:
@@
Figure 4: The architecture of LRM. The text encoder and UNet of diffusion models are utilized for preference prediction. The VAE encoder is only used in training.
where $gs\geq 1$ is a hyperparameter and $n_{m}$ denotes the dimension of $V_{mid}$. As $gs$ increases, $V_{mid}$ incorporates more features related to the text, leading to an enhanced focus on text-image alignment. When $gs=1$, the VFE module is not utilized. Next, $V_{enh}$ are concatenated with down-block features $V_{down}$ along the channel dimension and projected into the final visual features $V\in\mathbb{R}^{1\times n_{d}}$ by the visual projection layer. Finally, the preference score of $x_{t}$ and $p$ is computed as the dot product between textual and visual features after applying $l_{2}$ normalization:
@@
where $\tau$ denotes the temperature coefficient following CLIP.
Training Loss. LRM is trained on the public preference dataset Pick-a-Pic v1 (Kirstain et al., 2023), denoted as $D$, which consists of triples $(I^{w},I^{l},p)$, each containing a preference image pair and a corresponding prompt. As depicted in Fig. 3 (b), given uniformly sampled timestep $t\sim\mathcal{U}(0,T)$ where $T$ denotes the total denoising steps, we use the VAE encoder of the diffusion model to transform pixel-level images $I$ into latent images $x_{0}$, and then add $x_{0}$ with noise $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to simulate $x_{t}$ in the backward denoising process. Following the Bradley-Terry (BT) model (Bradley & Terry, 1952), the training loss of LRM is formulated as:
$L_{LRM}=-\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}(0,T),\\ (I^{w},I^{l},p)\in D\end{subarray}}\log\frac{e^{S(p,x_{t}^{w})}}{e^{S(p,x_{t}^{w})}+e^{S(p,x_{t}^{l})}},$ (11)
$x_{t}^{*}=\sqrt{\bar{\alpha}_{t}}*\text{VAE}_{\text{encoder}}(I^{*})+\sqrt{1-\bar{\alpha}_{t}}*\epsilon_{t}^{*},$ (12)
where $*\in\{w,l\}$. During training, the parameters of VAE are frozen to ensure that the latent space of the LRM remains stable, while the text encoder and UNet are trainable.
Figure 5: The distribution of preference score gaps in Pick-a-Pic.
@@ 4.2 Multi-Preference Consistent Filtering
Issue of Inconsistent Preferences. The training loss of LRM in Eqn. (11) involves the assumption that if $x^{w}_{0}$ is preferred over $x^{l}_{0}$, then after adding noise of equal intensity, the preference order remains unchanged, meaning $x^{w}_{t}$ continues to be preferred over $x^{l}_{t}$. However, this assumption breaks down when the winning image excels in one aspect but is inferior in other aspects. For example, if the winning image $x^{w}_{0}$ has better details but exhibits weaker text-image alignment compared to the losing image $x_{0}^{l}$, the advantage in detail may be diminished after introducing significant noise, making the noisy losing image $x_{t}^{l}$ potentially more preferred than $x_{t}^{w}$. Such situations frequently occur in the Pick-a-Pic (Kirstain et al., 2023) dataset. To demonstrate this, we employ the Aesthetic Score $S_{A}$ (Schuhmann et al., 2022) to evaluate the aesthetic quality while using CLIP Score $S_{C}$ (Radford et al., 2021) and VQAScore $S_{V}$ (Lin et al., 2025) to assess the text-image alignment. Let $G_{*}=S_{*}(I^{w})-S_{*}(I^{l}),*\in\{A,C,V\}$ denote the preference score gaps between winning and losing images. As illustrated in Fig. 5, nearly half of the winning images have Aesthetic Scores lower than those of losing images. Similarly, around 40% of the winning images exhibit lower CLIP Scores and VQAScores. When considering both Aesthetic and CLIP scores, over 70% of the winning images score lower than the losing images in at least one aspect.
Table 1: Explored data filter strategies of Pick-a-Pic.
@@ 4.3 Latent Preference Optimization
To utilize LRM for step-level preference optimization, we propose the Latent Preference Optimization (LPO) method.
Sampling and Training. As illustrated in Fig. 3 (c), at each timestep $t$, LPO samples a set of $x^{i}_{t},i=1,...,K$ from the same $x_{t+1}$. These noisy latent images are then directly fed into LRM to predict preference scores $S^{i}_{t},i=1,...,K$. The highest and lowest scores are normalized by the SoftMax function. If their score gap exceeds a predefined threshold $th_{t}$, the $x_{t}$ with the highest score is selected as $x_{t}^{w}$, while the one with the lowest score is designated as $x_{t}^{l}$, forming a qualified training sample $(x_{t+1},x_{t}^{w},x_{t}^{l})$. Finally, these samples are used to optimize the diffusion model, using the same loss with SPO, i.e. Eqn. (7). As a result, the entire process of LPO is performed within the latent space, eliminating the need to decode into pixel space, which significantly enhances the training efficiency of LPO.
Optimization Timesteps. Due to the inaccuracy of SPM at high-noise levels, SPO samples training data only at low to medium noise levels, specifically when $t\in[0,750]$. In contrast, owing to the noise-aware capability of LRM, LPO is able to collect training samples throughout the entire denoising process, covering $t\in[0,950]$. Experiments in Sec. 5.3 demonstrate that training samples collected at high noise levels play a crucial role in preference optimization.
@@ 5.1 Experimental Setup
The experiments are mainly conducted on SD1.5 (Rombach et al., 2022) and SDXL (Podell et al., 2023) without refiner. The LRM is first trained on Pick-a-Pic and then used to fine-tune diffusion models through LPO. Unless otherwise specified, we employ homogeneous optimization.
LRM Training. We denote the LRM based on SD1.5 and SDXL as LRM-1.5 and LRM-XL, respectively. They are trained on the filtered Pick-a-Pic v1 (Kirstain et al., 2023) as clarified in Sec. 4.2. The $gs$ in the VFE module is set to 7.5. More details are in Appendix A.
LPO Training. The same 4k prompts in SPO are used for the LPO training, randomly sampled from the training set of Pick-a-Pic v1. The DDIM scheduler (Song et al., 2021) with 20 inference steps is employed. We use all steps for sampling and training, i.e. $t\in[0,50,...,900,950]$. The dynamic threshold range $[th_{min},th_{max}]$ is set to $[0.35,0.5]$ for SD1.5 and $[0.45,0.6]$ for SDXL. The $\beta$ in Eqn. (7) is set to 500 and the $K$ in the sampling process is set to 4. Further details can be found in Appendix A.
Table 2: General and aesthetic preference scores on Pick-a-Pic validation unique set. denotes the metrics are copied from (Liang et al., 2024). Others are evaluated using the official model.
Method PickScore ImageReward HPSv2 HPSv2.1 Aesthetic
SD1.5
Original 20.56 0.0076 26.46 24.05 5.468
DDPO 21.06 0.0817 - 24.91 5.591
D3PO 20.76 -0.1235 - 23.97 5.527
Diff.-DPO 20.99 0.3020 27.03 25.54 5.595
SPO 21.22 0.1678 26.73 25.83 5.927
LPO 21.69 0.6588 27.64 27.86 5.945
SDXL
Original 21.65 0.4780 27.06 26.05 5.920
Diff.-DPO 22.22 0.8527 28.10 28.47 5.939
MaPO 21.89 0.7660 27.61 27.44 6.095
SPO 22.70 0.9951 28.42 31.15 6.343
LPO 22.86 1.2166 28.96 31.89 6.360
Table 3: Quantitative results on T2I-CompBench++ (Huang et al., 2023).
@@
SD1.5 Original (Rombach et al., 2022) 0.3783 0.3616 0.4172 0.1230 0.2967 0.4485 0.3104 0.2999
Diff.-DPO (Wallace et al., 2024) 0.4090 0.3664 0.4253 0.1336 0.3124 0.4543 0.3115 0.3042
SPO (Liang et al., 2024) 0.4112 0.4019 0.4044 0.1301 0.2909 0.4372 0.3008 0.2988
LPO 0.5042 0.4522 0.5259 0.1928 0.3562 0.4845 0.3110 0.3308
SDXL Original (Podell et al., 2023) 0.5833 0.4782 0.5211 0.1936 0.3319 0.4874 0.3137 0.3327
Diff.-DPO (Wallace et al., 2024) 0.6941 0.5311 0.6127 0.2153 0.3686 0.5304 0.3178 0.3525
MaPO (Hong et al., 2024) 0.6090 0.5043 0.5485 0.1964 0.3473 0.5015 0.3154 0.3229
SPO (Liang et al., 2024) 0.6410 0.4999 0.5551 0.2096 0.3629 0.4931 0.3098 0.3467
LPO 0.7351 0.5463 0.6606 0.2414 0.4075 0.5493 0.3152 0.3801
Table 4: Quantitative results on GenEval (Ghosh et al., 2023).
SD1.5 Original 97.50 37.12 34.69 75.53 3.75 6.75 42.56
Diff.-DPO 98.44 38.38 36.25 77.93 4.50 7.25 43.79
SPO 95.00 33.84 32.50 69.95 4.25 7.25 40.46
LPO 97.81 54.80 40.94 79.52 7.00 10.25 48.39
SDXL Original 93.75 63.38 30.94 80.05 9.25 19.00 49.40
Diff.-DPO 99.06 76.52 45.00 88.83 11.50 25.75 57.78
MaPO 95.63 68.94 32.19 83.51 11.50 17.75 51.59
SPO 94.38 69.44 31.88 81.65 10.25 15.50 50.52
LPO 99.69 81.57 43.75 89.10 14.00 27.50 59.27
Table 5: Comparisons of training speed.
@@
LPO 52 40 92
Table 6: Heterogeneous optimization based on LRM-SD1.5. P-S and I-R denote the PickScore and ImageReward metrics.
Model Method Aesthetic GenEval P-S I-R HPSv2 HPSv2.1
SD2.1 Original 5.673 48.59 20.92 0.3063 27.05 25.49
(Same VAE) LPO 5.969 56.01 21.76 0.7978 28.05 28.61
SDXL Original 5.920 49.40 21.65 0.4780 27.06 26.05
(Diff. VAE) LPO 5.953 40.85 20.82 0.3919 27.10 26.69
Baseline Methods. We compare LPO with DDPO (Black et al., 2024), D3PO (Yang et al., 2024), Diffusion-DPO (Wallace et al., 2024), MaPO (Hong et al., 2024), and SPO (Liang et al., 2024). These methods are trained on similar datasets, such as Pick-a-Pic v1 and v2, to ensure a fair comparison. Details are provided in Appendix A.
Evaluation Protocol. We evaluate various diffusion models across three dimensions: general preference, aesthetic preference, and text-image alignment. The PickScore (Kirstain et al., 2023), HPSv2 (Wu et al., 2023), HPSv2.1 (Wu et al., 2023), and ImageReward (Xu et al., 2023) are utilized to assess the general preference. The aesthetic preference is evaluated using the Aesthetic Score (Schuhmann et al., 2022). Consistent with (Liang et al., 2024), both general and aesthetic preferences are assessed on the validation unique split of Pick-a-Pic v1, which has 500 different prompts. For text-image alignment, we employ the GenEval (Ghosh et al., 2023) and T2I-CompBench++ (Huang et al., 2023) metrics. All images are generated using the DDIM scheduler with 20 steps. Additionally, to assess the correlations between the LRM and aesthetics as well as text-image alignment, we propose two corresponding metrics. Specifically, we calculate the score gaps $G_{*},*\in\{A,C,L\}$ between winning and losing images, where $A$, $C$, $L$ represent Aesthetic, CLIP, and LRM. For LRM, the score is taken at $t=0$. Then the Pearson Correlation Coefficient (Pearson, 1896) between $G_{L}$ and $G_{A}$ is referred to as Aes-Corr while that between $G_{L}$ and $G_{C}$ is termed CLIP-Corr. They are evaluated on the validation unique and test unique splits of Pick-a-Pic v1.

### v3

@@ 4.1 Latent Reward Model
Figure 3: The architecture of LRM. The VAE encoder is only used during training.
Architecture of LRM. LRM leverages features of the U-Net and text encoder in the diffusion model for preference prediction, as depicted in Fig. 3. Specifically, the textual features $f_{p}\in\mathbb{R}^{n_{l}\times n_{p}}$ are extracted from the prompt $p$ by the text encoder, where $n_{l}$ and $n_{p}$ denote the number of tokens and the dimension of textual features, respectively. Following CLIP [9], we use the last token $f_{eos}\in\mathbb{R}^{1\times n_{p}}$ to represent the entire prompt and incorporate a text projection layer to obtain the final textual features $T\in\mathbb{R}^{1\times n_{d}}$, where $n_{d}$ represents the final dimension of textual and visual features. Each noisy latent image $x_{t}$ is passed through the U-Net to interact with textual features $f_{p}$. The visual features of U-Net are averagely pooled along the spatial dimension, resulting in the multiscale down-block features $V_{down}$ and middle-block features $V_{mid}$ as follows
$V_{down},V_{mid}=\text{AvgPool}(\text{U-Net}(x_{t},f_{p})),\quad V_{down}=\{V_{d_{i}},i=1,\ldots,L\},$ (7)
where $L$ is the number of down blocks and $V_{d_{i}}$ represents features of the $i$-th down blocks. However, it is observed that these features lack sufficient correlations with the textual features. Inspired by the Classifier-Free Guidance [36], we propose the Visual Feature Enhancement (VFE) module to enhance correlations between visual and textual features. It first extracts middle-block features $V_{mid\_ucond}$ without textual information, following a process similar to extracting $V_{mid}$, but with a null prompt. The enhanced visual features $V_{enh}\in\mathbb{R}^{1\times n_{m}}$ are then computed as follows:
@@
where $gs\geq 1$ is a hyperparameter and $n_{m}$ is the dimension of $V_{mid}$. A larger $gs$ value injects more text-related features into $V_{mid}$, thus enhancing its focus on text-image alignment. When $gs=1$, the VFE module is disabled. Next, $V_{enh}$ is concatenated with down-block features $V_{down}$ along the channel dimension and projected into the final visual features $V\in\mathbb{R}^{1\times n_{d}}$ via a visual projection layer. Finally, the preference score of $x_{t}$ and $p$ is the dot product between textual and visual features:
@@
where $\tau$ is a temperature coefficient following CLIP [9] and $l_{2}$ denotes the L2 Norm.
Figure 4: The training pipelines of PRM (a) and LRM (b), and sampling pipelines of SPO (c) and LPO (d). Compared to PRM, LRM can directly process latent images $x_{t}$ at various noise levels without requiring $\hat{x}_{0,t}$ prediction and VAE decoding. Therefore, LPO can perform sampling entirely within the noisy latent space throughout the full denoising process, covering $t\in[0,950]$.
Training Loss. LRM is trained on the public preference dataset Pick-a-Pic v1 [16], denoted as $D$, which consists of triples, each containing a preferred image $I^{w}$, a less preferred image $I^{l}$, and the corresponding prompt $p$. As depicted in Fig. 4 (b), given uniformly sampled timestep $t\sim\mathcal{U}(0,T)$ where $T$ denotes the total denoising steps, the pixel-level image $I$ is encoded into the latent image $x_{0}$ using the VAE encoder. Noise $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is then added to simulate $x_{t}$ in the backward denoising process. Following the Bradley-Terry (BT) model [37], the training loss of LRM is formulated as:
$L_{LRM}=-\mathbb{E}_{t\sim\mathcal{U}(0,T),(I^{w},I^{l},p)\in D}\log\frac{exp(S(p,x_{t}^{w}))}{exp(S(p,x_{t}^{w}))+exp(S(p,x_{t}^{l}))},\$ (10)
$x_{t}^{*}=\sqrt{\bar{\alpha}_{t}}*\text{VAE}_{\text{encoder}}(I^{*})+\sqrt{1-\bar{\alpha}_{t}}*\epsilon_{t}^{*},\quad*\in\{w,l\}.$ (11)
During training, the parameters of VAE are frozen to ensure that the latent space of the LRM remains stable, while the text encoder and U-Net are trainable.
@@ 4.2 Multi-Preference Consistent Filtering
Figure 5: The distribution of preference consistency between human labels and reward scores in Pick-a-Pic.
Issue of Inconsistent Preferences. The training loss of LRM in Eqn. (10) involves the assumption that if $x^{w}_{0}$ is preferred over $x^{l}_{0}$, then after adding noise of equal intensity, the preference order remains unchanged, meaning $x^{w}_{t}$ continues to be preferred over $x^{l}_{t}$. However, this assumption breaks down when the winning image excels in one aspect but is inferior in other aspects. For example, if the winning image $x^{w}_{0}$ has better details but exhibits weaker text-image alignment compared to the losing image $x_{0}^{l}$, the advantage in detail may be diminished after introducing significant noise, making the noisy losing image $x_{t}^{l}$ potentially more preferred than $x_{t}^{w}$. Such situations frequently occur in the Pick-a-Pic v1 [16] dataset. To demonstrate this, we employ the Aesthetic Score $S_{A}$ [38] to evaluate the aesthetic quality while using CLIP Score $S_{C}$ [9] and VQAScore $S_{V}$ [39] to assess the text-image alignment. Let $G_{*}=S_{*}(I^{w})-S_{*}(I^{l}),*\in\{A,C,V\}$ denote the preference score gaps between winning and losing images. As illustrated in Fig. 5, nearly half of the winning images have Aesthetic Scores lower than those of losing images. Similarly, around 40% of the winning images exhibit lower CLIP Scores and VQAScores. When considering both Aesthetic and CLIP scores, over 70% of the winning images score lower than the losing images in at least one aspect.
Table 1: Data filter strategies for Pick-a-Pic.
@@ 4.3 Latent Preference Optimization
To utilize LRM for step-level preference optimization, we propose the Latent Preference Optimization (LPO) method. We also explore the application of LRM on GRPO [15] in Appendix B.
Sampling and Training. As illustrated in Fig. 4 (d), at each timestep $t$, LPO samples a set of $x^{i}_{t},i=1,...,K$ from the same $x_{t+1}$. These noisy latent images are then directly fed into LRM to predict preference scores $S^{i}_{t},i=1,...,K$. The highest and lowest scores are normalized by the SoftMax function. If their score gap exceeds a predefined threshold $th_{t}$, the $x_{t}$ with the highest score is selected as $x_{t}^{w}$, while the one with the lowest score is designated as $x_{t}^{l}$, forming a qualified training sample $(x_{t+1},x_{t}^{w},x_{t}^{l})$. Finally, these samples are used to optimize the diffusion model using the loss in Eqn. (6). As a result, the entire process of LPO is conducted within the noisy latent space, eliminating the need for $\boldsymbol{\hat{x}_{0,t}}$ prediction and VAE decoding, which significantly enhances the training effectiveness and efficiency of LPO.
Optimization Timesteps. Due to the inaccuracy of SPM at high-noise levels, SPO samples training data only at low to medium noise levels, specifically when $t\in[0,750]$, as shown in Fig. 4 (c). In contrast, owing to LRM’s noise-aware ability, LPO can perform sampling and training throughout the entire denoising process, covering $t\in[0,950]$. Experiments in Sec. 5.3 and Appendix B demonstrate that optimization at high noise levels plays a crucial role in preference optimization.
@@ 5.1 Experimental Setup
The experiments are mainly conducted on SD1.5 [14] and SDXL [11] without refiner. The LRM is first trained on Pick-a-Pic and then used to fine-tune diffusion models through LPO. If not specified, we employ homogeneous optimization. Ablation experiments are conducted on SD1.5.
LRM and LPO Training. We denote the LRM based on SD1.5 and SDXL as LRM-1.5 and LRM-XL, respectively. They are trained on the filtered Pick-a-Pic v1 [16] as clarified in Sec. 4.2. The $gs$ in the VFE module is set to 7.5. The same 4k prompts in SPO are used for the LPO training, randomly sampled from the training set of Pick-a-Pic v1. The DDIM scheduler [35] with 20 inference steps is employed. We use all steps for sampling and training, i.e. $t\in[0,50,...,900,950]$. The dynamic threshold range $[th_{min},th_{max}]$ is set to $[0.35,0.5]$ for SD1.5 and $[0.45,0.6]$ for SDXL. The $\beta$ in Eqn. (6) is set to 500 and the $K$ in the sampling process is set to 4. More details are in Appendix A.
Baseline Methods. We compare LPO with DDPO [6], D3PO [7], Diffusion-DPO [4], MaPO [29], SPO [8], and SePPO [30]. These methods are trained on Pick-a-Pic, which ensures a relatively fair comparison. We also include InterComp [31] for reference, which uses a higher-quality internal dataset. More details are provided in Appendix A.
Evaluation Protocol. We evaluate various diffusion models across three dimensions: general preference, aesthetic preference, and text-image alignment. The PickScore [16], HPSv2 [17], HPSv2.1 [17], and ImageReward [18] are utilized to assess the general preference. The aesthetic preference is evaluated using the Aesthetic Score [38]. Following [8], both general and aesthetic preferences are assessed on the validation unique split of Pick-a-Pic v1, which has 500 different prompts. For text-image alignment, we employ the GenEval [40] and T2I-CompBench++ [41] metrics. For detailed evaluation practice, please refer to Appendix A. Additionally, we propose two metrics to assess LRM’s correlations with aesthetics and text-image alignment. Specifically, we calculate the score gaps $G_{*},*\in\{A,C,L\}$ between winning and losing images, where $A$, $C$, $L$ represent Aesthetic, CLIP, and LRM. For LRM, the score is taken at $t=0$. Then the Pearson Correlation Coefficient [42] between $G_{L}$ and $G_{A}$ is referred to as Aes-Corr while that between $G_{L}$ and $G_{C}$ is termed CLIP-Corr. They are evaluated on the validation unique and test unique splits of Pick-a-Pic v1.
Table 2: General and aesthetic preference scores on Pick-a-Pic validation unique set, along with GenEval [40] scores. denotes the metrics are copied from [8]. Others are evaluated using the official model. If not specified, we use 20 inference steps. PaP denotes the Pick-a-Pic [16] dataset. The complete results on GenEval are provided in Tab. 26 and Tab. 26.
Model Method Dataset PickScore ImageReward HPSv2 HPSv2.1 Aesthetic GenEval (20 Step) GenEval (50 Step)
SD1.5 [14] Original - 20.56 0.0076 26.46 24.05 5.468 42.56 42.53
DDPO [6] PaP v2 21.06 0.0817 - 24.91 5.591 - -
D3PO [7] PaP v1 20.76 -0.1235 - 23.97 5.527 - -
Diff.-DPO [4] PaP v2 20.99 0.3020 27.03 25.54 5.595 43.79 45.13
SPO [8] PaP v1 21.43 0.1712 - 26.45 5.887 - -
SPO [8] PaP v1 21.22 0.1678 26.73 25.83 5.927 40.46 41.53
SePPO [30] PaP v2 21.25 0.5077 27.56 27.34 5.766 44.97 44.46
LPO (ours) PaP v1 21.69 0.6588 27.64 27.86 5.945 48.39 48.77
SDXL [11] Original - 21.65 0.4780 27.06 26.05 5.920 49.40 52.29
Diff.-DPO [4] PaP v2 22.22 0.8527 28.10 28.47 5.939 57.78 58.91
MaPO [29] PaP v2 21.89 0.7660 27.61 27.44 6.095 51.59 52.80
SPO [8] PaP v1 23.06 1.0803 - 31.80 6.364 - 55.20
SPO [8] PaP v1 22.70 0.9951 28.42 31.15 6.343 50.52 52.75
InterComp [31] Internal 22.63 1.2728 29.08 31.52 6.016 59.24 59.65
LPO (ours) PaP v1 22.86 1.2166 28.96 31.89 6.360 59.27 59.85
Table 3: Quantitative results on T2I-CompBench++ [41] with 20 inference steps.
@@
SD1.5 [14] Original [14] 0.3783 0.3616 0.4172 0.1230 0.2967 0.4485 0.3104 0.2999
Diff.-DPO [4] 0.4090 0.3664 0.4253 0.1336 0.3124 0.4543 0.3115 0.3042
SPO [8] 0.4112 0.4019 0.4044 0.1301 0.2909 0.4372 0.3008 0.2988
SePPO [30] 0.4265 0.3747 0.4170 0.1504 0.3285 0.4568 0.3109 0.3076
LPO (ours) 0.5042 0.4522 0.5259 0.1928 0.3562 0.4845 0.3110 0.3308
SDXL [11] Original [11] 0.5833 0.4782 0.5211 0.1936 0.3319 0.4874 0.3137 0.3327
Diff.-DPO [4] 0.6941 0.5311 0.6127 0.2153 0.3686 0.5304 0.3178 0.3525
MaPO [29] 0.6090 0.5043 0.5485 0.1964 0.3473 0.5015 0.3154 0.3229
SPO [8] 0.6410 0.4999 0.5551 0.2096 0.3629 0.4931 0.3098 0.3467
InterComp [31] 0.7218 0.5335 0.6290 0.2406 0.3929 0.5395 0.3212 0.3659
LPO (ours) 0.7351 0.5463 0.6606 0.2414 0.4075 0.5493 0.3152 0.3801
Table 4: Heterogeneous optimization based on LRM-SD1.5.
Model Method Aesthetic GenEval P-S I-R HPSv2 HPSv2.1
SD2.1 [14] Original 5.673 48.59 20.92 0.3063 27.05 25.49
(Same VAE) LPO (ours) 5.969 56.01 21.76 0.7978 28.05 28.61
SDXL [11] Original 5.920 49.40 21.65 0.4780 27.06 26.05
(Diff. VAE) LPO (ours) 5.953 40.85 20.82 0.3919 27.10 26.69
Table 5: Time of each sampling step on a single A100.
Reward Model Denoising Step ($x_{t+1}\rightarrow x_{t}$) $\hat{x}_{0,t}$ Prediction ($x_{t}\rightarrow\hat{x}_{0,t}$) VAE Decoding ($\hat{x}_{0,t}\rightarrow I_{t}$) Reward Prediction Total $\downarrow$ (K=4)
SPM (PRM) 0.019s $K\times$ 0.019s $K\times$ 0.031s $K\times$ 0.006s 0.243s
LRM (ours) 0.019s 0 0 $K\times$ 0.005s 0.039s
Table 6: Comparisons of training speed between different methods.
@@
LPO (ours) 52 40 92